# Continuous fields of interval algebras

### v1

@@ B. The uniqueness
We aim to extend the uniqueness part of Section 3 to morphisms between continuous fields whose fibers fall into the class of separable unital $\operatorname{AI}$-algebras. As for the existence part, restrictions on the base space and the fibers will appear. The technical key is inspired from a ‘gluing process’ developed in [9, Lemma 3.4], used to extend from local-to-global, the classification of $\operatorname{AF}$-algebras by means of the $\mathrm{K}_{0}$-group to certain continuous fields whose fibers fall into the class of $\operatorname{AF}$-algebras. Nevertheless, their argumentation has to be refined for our context, which generally has $\mathrm{K}_{1}$-obstructions. Our gluing process is based on the following result.
[…]
[5, Theorem 2.3] Let $A$ be a unital interval algebra. Let $F\subseteq A$ be a finite set containing the canonical generators of $A$ and let $\epsilon>0$. Then there exists $\delta(F,\epsilon)>0$ satisfying the following.
@@
For any continuous field morphisms $\phi,\psi:\mathcal{A}\to\mathcal{B}$ and any $w_{0},w_{1}\in U(B)$ such that
$(w_{0}\phi w_{0}^{*})_{y}\simeq_{(F_{y},\delta)}\psi_{y}\simeq_{(F_{y},\delta)}(w_{1}\phi w_{1}^{*})_{y}\quad\forall y\in Y$
there exists $w\in U(\mathcal{B})$ with $w(y_{0})=w_{0}(y_{0})$ and $w(y_{1})=w_{1}(y_{1})$, such that
$(w\phi w^{*})_{y}\simeq_{(F_{y},\epsilon)}\psi_{y}\quad\forall y\in Y$
@@
We first define the unitary element $u:=w_{1}^{*}w_{0}$ of $\mathcal{B}$. Combining the equalities of ()($\ast$)( ∗ ) we deduce that $(u\phi u^{*})_{y}\simeq_{(F_{y},2\delta)}\psi_{y}$ for all $y\in Y$, and a fortiori at the fiber $y_{0}$. Now choose $\delta>0$ small enough such that $2\delta<\min(\delta(F_{y_{0}},\epsilon/3),\epsilon/3)$ where $\delta(F_{y_{0}},\epsilon/3)$ is obtained from  B. As a consequence, there exists a path $\gamma_{0}$ in $U(B_{y_{0}})$ joining $u(y_{0})$ to $1_{B_{y_{0}}}$ such that
[…]
Since $\mathcal{B}\twoheadrightarrow B_{y_{0}}$ is a surjective unital *-homomorphism, it is known that elements of $U_{0}(B_{y_{0}})$ lift to elements of $U_{0}(\mathcal{B})$. (See e.g. [16, Corollary 4.3.3].) Therefore, there exists a path $\gamma$ in $U_{0}(\mathcal{B})$ joining $u$ to $1_{\mathcal{B}}$ whose projection to the fiber $B_{y_{0}}$ satisfies that $\gamma_{y_{0}}=\gamma_{0}$. By the continuity of the norm and the finiteness of $F$, we can find some $y_{\epsilon}\in Y$ such that
[…]
@@
$\begin{array}[]{ll}\widetilde{u}\colon[0,1]\longrightarrow\bigsqcup_{x}\mathcal{B}_{x}\\ \hskip 28.45274ptx\longmapsto\Biggl{\{}\begin{array}[]{ll}u(x)&\text{ if}\quad 0\leq x\leq y_{0}\\ \gamma_{x}(h(x))&\text{ if}\quad y_{0}<x\leq y_{\epsilon}\\ 1_{B_{y}}&\text{ if}\quad y_{\epsilon}<x\leq 1\end{array}\end{array}$
@@
$\leq 2\epsilon/3+\delta$
$\leq\epsilon$
@@
[…]
Let $([0,1],\{A_{x}\}_{x},\mathcal{A})$ and $([0,1],\{B_{x}\}_{x},\mathcal{B})$ be continuous fields of separable unital $\operatorname{AI}$-algebras. Assume moreover that the fibers of $\mathcal{A}$ are interval algebras. Let $\phi,\psi:\mathcal{A}\to\mathcal{B}$ be continuous field morphisms. For any finite set $F\subseteq\mathcal{A}$ and any $\epsilon>0$, there exists a finite set $G\subseteq\operatorname{Cu}(\mathcal{A})$ satisfying the following.
@@
Let $x\in X$. We know by Section 3 that for any $\delta>0$, there exists a finite set $G_{x}\subseteq\operatorname{Cu}(A_{x})$ satisfying the following: If $\operatorname{Cu}(\phi_{x})\simeq_{G_{x}}\operatorname{Cu}(\psi_{x})$, then there exists a unitary element $w_{x}\in B_{x}$ such that $w_{x}\phi_{x}w_{x}^{*}\simeq_{(F_{x},\delta/2)}\psi_{x}$. Arguing similarly as in the proof of  B, we deduce that for any $\delta>0$ we can find a unitary element $w\in U_{0}(\mathcal{B})$ and an open neighborhood $V$ of $x$ such that
$(\ast_{x})\quad\operatorname{Cu}(\phi_{x})\simeq_{G_{x}}\operatorname{Cu}(\psi_{x})\Rightarrow\exists w\in U_{0}(\mathcal{B})\mid(w\phi w^{*})_{v}\simeq_{(F_{v},\delta)}\psi_{v}\quad\forall v\in V.$
Moreover, there exists a finite set $G\subseteq\operatorname{Cu}(\mathcal{A})$ whose projection on the fiber $\operatorname{Cu}(A_{x})$ is $G_{x}$.
It follows from the compactness and the dimension of $[0,1]$ that there exist finitely many points $x_{1},\dots,x_{l}$ together with open neighborhoods $V_{1},\dots,V_{l}$, unitary elements $w_{1},\dots,w_{l}$ of $\mathcal{B}$ and finite sets $G^{(1)},\dots,G^{(l)}$ of $\operatorname{Cu}(\mathcal{A})$, such that $\bigcup_{1}^{l}V_{i}$ is a brick cover of $[0,1]$ (i.e., a cover such that $V_{i}\cap V_{j}=\emptyset$ whenever $|i-j|>1$) and yielding the following implications. For any $1\leq i\leq l$
$(\ast_{i})\quad\operatorname{Cu}(\phi_{x_{i}})\simeq_{G^{(i)}_{x_{i}}}\operatorname{Cu}(\psi_{x_{i}})\Rightarrow(w_{i}\phi w_{i}^{*})_{v}\simeq_{(F_{v},\delta)}\psi_{v}\quad\forall v\in V_{i}$
Subsequently, we recursively modify the unitary elements $\{\omega_{i}\}_{1}^{l}$ in order to construct the global unitary we are looking for. More specifically, for each $11\leq i\leq l-$ 1, we fix a subinterval $Y_{i}:=[y_{i,0},y_{i,1}]\subsetneq V_{i}\cap V_{i+1}$ and $\delta>0$ small enough (in particular, we choose $\delta<\epsilon$) in order to be able to invoke  B and obtain a unitary element $\widetilde{w}_{i}$ of $\mathcal{B}$ satisfying the following
$\biggl{\{}\begin{array}[]{ll}(\widetilde{w}_{i}\phi\widetilde{w}_{i}^{*})_{y}\simeq_{(F_{y},\epsilon)}\psi_{y}\quad\forall y\in Y_{i}\\ \widetilde{w}_{i}(y_{i,0})=w_{i}(y_{i,0})\text{ and }\widetilde{w}_{i}(y_{i,1})=w_{i+1}(y_{i,1})\end{array}$
@@
$\begin{array}[]{ll}w\colon[0,1]\longrightarrow\bigsqcup_{x}\mathcal{B}_{x}\\ \hskip 28.45274ptx\longmapsto\biggl{\{}\begin{array}[]{ll}w_{i}(x)&\text{ if}\quad x\in V_{i}\setminus Y_{i}\\ \widetilde{w}_{i}(x)&\text{ if}\quad x\in Y_{i}\end{array}\end{array}$
Let us set $G:=\bigcup_{1}^{l}G^{(i)}$. From construction, we know that $\operatorname{Cu}(\phi_{x})\simeq_{G_{x}}\operatorname{Cu}(\psi_{x})$ for any $x\in[0,1]$. A fortiori, $\operatorname{Cu}(\phi_{x_{i}})\simeq_{G^{(i)}_{x_{i}}}\operatorname{Cu}(\psi_{x_{i}})$ for any $1\leq i\leq l$, from which we conclude that $(w\phi w^{*})_{x}\simeq_{(F_{x},\epsilon)}\psi_{x}$ for any $x\in\bigcup_{1}^{l}V_{i}=[0,1].$
[…]
@@
$\|w_{n}\phi w_{n}^{*}(a)-\psi(a)\|_{\mathcal{A}}\underset{n\infty}{\longrightarrow}0\quad\forall a\in\mathcal{A}$
@@ C. The classification
Firstly, we say that two continuous field morphisms $\phi,\psi\colon(X,\{A_{x}\}_{x},\mathcal{A})\rightarrow(X,\{B_{x}\}_{x},\mathcal{B})$ are approximately unitarily equivalent if there exists a sequence $(w_{n})_{n}$ of unitaries of $\mathcal{B}$ such that $w_{n}\phi w_{n}^{*}\underset{n\infty}{\rightarrow}\psi$ in the point-norm topology. Equivalently, if they are approximately unitarily equivalent as *-homomorphisms.
Secondly, we say that a pair $[(X,\{A_{x}\}_{x},\mathcal{A}),(X,\{B_{x}\}_{x},\mathcal{B})]$ of continuous fields satisfies the classification property, termed (CP) for short, if for any scaled $\operatorname{Cu}$-morphism $\alpha\colon\operatorname{Cu}(\mathcal{A})\rightarrow\operatorname{Cu}(\mathcal{B})$ descending to the fibers, there exists a continuous field morphism $\phi\colon\mathcal{A}\rightarrow\mathcal{B}$, unique up to approximate unitary equivalence, such that $\operatorname{Cu}(\phi)=\alpha$. We may write ‘$[\mathcal{A},\mathcal{B}]$ has (CP)’ when the context is clear.
@@
[…]
Let $(X,\{(A_{i})_{x}\}_{x},\mathcal{A}_{i})_{i}$ be a family of continuous fields indexed over an upward-directed set. Consider any inductive system $((X,\{(A_{i})_{x}\}_{x},\mathcal{A}_{i}),\phi_{ij})_{i}$ in $\mathcal{F}_{C^{*}}$. We have the following.
(i) $\mathcal{F}_{C^{*}}\text{-}\lim(\mathcal{A}_{i},\phi_{ij})\simeq C^{*}\text{-}\lim(\mathcal{A}_{i},\phi_{ij})$.
(ii) $[\mathcal{F}_{C^{*}}\text{-}\lim(\mathcal{A}_{i},\phi_{ij}),\mathcal{B}]$ has (CP) whenever $[\mathcal{A}_{i},\mathcal{B}]$ has (CP) for all $i\in I$.
@@
(i) Consider the inductive system in the category $C^{*}$ and denote by $(\mathcal{A},\phi_{i\infty})$ its $C^{*}$-limit. For any $x\in X$, consider the inductive system obtained by the projection of the objects and morphisms at the fiber $x$. We obtain another inductive system in $C^{*}$ and we denote its limit by $(A_{x},(\phi_{i\infty})_{x})$. A standard one-sided intertwining argument shows that there exists a unique *-homomorphism $\pi_{x}\colon\mathcal{A}\rightarrow A_{x}$ compatible with the squares of the intertwining. A fortiori, $(X,\{A_{x}\}_{x},\mathcal{A})$ is a well-defined continuous field over $X$ and $\phi_{i\infty}$ is a continuous field morphism for any $i\in I$. It is now readily checked that $(\mathcal{A},\phi_{i\infty})$ satisfies the universal property of inductive limits in the category $\mathcal{F}_{C^{*}}$.
(ii) This is done similarly as in the proof of [8, Proposition 5.2 (iv)]. ∎
In what follows, $\mathcal{C}$ denotes the class of continuous fields that are realized as $\underset{\rightarrow}{\lim}(C[0,1]\otimes A_{i},\phi_{ii+1})_{i\in\mathbb{N}}$, where $A_{i}$ are unital interval algebras and $\phi_{ii+1}$ are unital continuous fields morphisms.
@@
Combine  A with  B and the fact that $[\mathcal{A},\mathcal{B}]$ has (CP). ∎
[…]
Let $\mathcal{A},\mathcal{B}\in\mathcal{C}$. Then any $\operatorname{Cu}$-isomorphism $\alpha\colon\operatorname{Cu}(A)\simeq\operatorname{Cu}(B)$ descending to the fibers and such that $\alpha([1_{A}])=[1_{B}]$, lifts to a continuous field isomorphism $\mathcal{A}\simeq\mathcal{B}$ which is unique up to approximate unitary equivalence.

### v2

@@ B. The uniqueness
We aim to extend the uniqueness part of Section 3 to morphisms between continuous fields whose fibers fall into the class of separable unital $\operatorname{AI}$-algebras. As for the existence part, restrictions on the base space and the fibers will appear. The technical key is inspired from a ‘gluing process’ developed in [14, Lemma 3.4], used to extend from local-to-global, the classification of $\operatorname{AF}$-algebras by means of the $\mathrm{K}_{0}$-group to certain continuous fields whose fibers fall into the class of $\operatorname{AF}$-algebras. Nevertheless, their argumentation has to be refined for our context, which generally has $\mathrm{K}_{1}$-obstructions. Our gluing process is based on the following result.
Theorem 3.18.
[6, Theorem 2.3] Let $A$ be a unital interval algebra. Let $F\subseteq A$ be a finite set containing the canonical generators of $A$ and let $\epsilon>0$. Then there exists $\delta(F,\epsilon)>0$ satisfying the following.
@@
For any continuous field morphisms $\phi,\psi:\mathcal{A}\to\mathcal{B}$ and any $w_{0},w_{1}\in U(\mathcal{B})$ such that
$(\ast)\quad\quad\quad(w_{0}\phi w_{0}^{*})_{y}\simeq_{(F_{y},\delta)}\psi_{y}\simeq_{(F_{y},\delta)}(w_{1}\phi w_{1}^{*})_{y},\quad\forall y\in Y,$
then there exists $w\in U(\mathcal{B})$ with $w(y_{0})=w_{0}(y_{0})$ and $w(y_{1})=w_{1}(y_{1})$, such that
$(w\phi w^{*})_{y}\simeq_{(F_{y},\epsilon)}\psi_{y},\quad\forall y\in Y.$
@@
We first consider the unitary element $u:=w_{1}^{*}w_{0}$ of $\mathcal{B}$. Combining the comparisons of $(\ast)$, we deduce that $(u\phi u^{*})_{y}\simeq_{(F_{y},2\delta)}\phi_{y}$ for all $y\in Y$, and a fortiori at the fiber $y_{0}$. Now choose $\delta>0$ small enough such that $2\delta<\min(\delta(F_{y_{0}},\epsilon/3),\epsilon/3)$ where $\delta(F_{y_{0}},\epsilon/3)$ is obtained from  B. As a consequence, there exists a path $\gamma_{0}$ in $U(B_{y_{0}})$ joining $u(y_{0})$ to $1_{B_{y_{0}}}$ such that
$\gamma_{0}(s)\phi_{y_{0}}\gamma_{0}(s)^{*}\simeq_{(F_{y_{0}},\epsilon/3)}\phi_{y_{0}}$
for any $0\leq s\leq 1$.
Since $\mathcal{B}\twoheadrightarrow B_{y_{0}}$ is a surjective unital *-homomorphism, it is known that elements of $U_{0}(B_{y_{0}})$ lift to elements of $U_{0}(\mathcal{B})$. (See e.g. [23, Corollary 4.3.3].) Therefore, there exists a path $\gamma$ in $U_{0}(\mathcal{B})$ joining $u$ to $1_{\mathcal{B}}$ whose projection to the fiber $B_{y_{0}}$ satisfies that $\gamma_{y_{0}}=\gamma_{0}$. By the continuity of the norm and the finiteness of $F$, we can find some $y_{\epsilon}\in Y$ such that
$\gamma_{y}(s)\phi_{y}\gamma_{y}(s)^{*}\simeq_{(F_{y},2\epsilon/3)}\phi_{y}$
for any $0\leq s\leq 1$ and any $y\in[y_{0},y_{\epsilon}]$.
@@
$\begin{array}[]{ll}\widetilde{u}\colon[0,1]\longrightarrow\bigsqcup_{x}B_{x}\\ \hskip 28.45274ptx\longmapsto\Biggl\{\begin{array}[]{ll}u(x)&\text{ if}\quad 0\leq x\leq y_{0}\\ \gamma_{x}(h(x))&\text{ if}\quad y_{0}<x\leq y_{\epsilon}\\ 1_{B_{y}}&\text{ if}\quad y_{\epsilon}<x\leq 1\end{array}\end{array}$
@@
$\leq 2\epsilon/3+\delta\leq\epsilon$
@@
Theorem 3.21.
Let $([0,1],\{A_{x}\}_{x},\mathcal{A})$ and $([0,1],\{B_{x}\}_{x},\mathcal{B})$ be continuous fields of separable unital $\operatorname{AI}$-algebras. Assume moreover that the fibers of $\mathcal{A}$ are interval algebras.
Let $\phi,\psi:\mathcal{A}\to\mathcal{B}$ be continuous field morphisms. For any finite set $F\subseteq\mathcal{A}$ and any $\epsilon>0$, there exists a finite set $G\subseteq\operatorname{Cu}(\mathcal{A})$ satisfying the following.
@@
Let $x\in X$ and let $\delta>0$. By the uniform continuity of the classification of Section 3, we can find a subset $H\subseteq\operatorname{Cu}(\mathcal{A})$ such that $H_{x}$ is $(F_{x},\delta)$-admissible for $A_{x}$. (See  A.) Using lifting of unitaries arguments similarly as in the proof of  B, there exists a unitary element $w\in U_{0}(\mathcal{B})$ such that $(w\phi w^{*})_{x}\simeq_{(F_{x},\delta)}\psi_{x}$, whenever $\operatorname{Cu}(\phi_{x})\simeq_{H_{x}}\operatorname{Cu}(\psi_{x})$. Further, by Lemma 3.6, we can find an open an open neighborhood $V$ of $x$ such that $(w\phi w^{*})_{v}\simeq_{(F_{v},\delta)}\psi_{v}$ for any $v\in V$, whenever $\operatorname{Cu}(\phi_{x})\simeq_{H_{x}}\operatorname{Cu}(\psi_{x})$.
Subsequently, it follows from the compactness and the dimension of $[0,1]$ that there exist finitely many points $x_{1},\dots,x_{l}$ together with open neighborhoods $V_{1},\dots,V_{l}$, unitary elements $w_{1},\dots,w_{l}$ of $\mathcal{B}$ and finite sets $G^{(1)},\dots,G^{(l)}$ of $\operatorname{Cu}(\mathcal{A})$, satisfying the followings.
$\bullet$ $\bigcup_{1}^{l}V_{i}$ is a cover of $[0,1]$ such that $V_{i}\cap V_{j}=\emptyset$ whenever |ij|>1|i-j|>1.
$\bullet$ For any $1\leq i\leq l$ and any $v\in V_{i}$, we have that $(w_{i}\phi w_{i}^{*})_{v}\simeq_{(F_{v},\delta)}\psi_{v}$, whenever $\operatorname{Cu}(\phi_{x_{i}})\simeq_{G^{(i)}_{x_{i}}}\operatorname{Cu}(\psi_{x_{i}})$.
Set $G:=\bigcup_{1}^{l}G^{(i)}$ and assume that $\operatorname{Cu}(\phi_{x})\simeq_{G_{x}}\operatorname{Cu}(\psi_{x})$ for all $x$. In particular, for any $1\leq i\leq l$, we have $\operatorname{Cu}(\phi_{x_{i}})\simeq_{G^{(i)}_{x_{i}}}\operatorname{Cu}(\psi_{x_{i}})$ and hence, $(w_{i}\phi w_{i}^{*})_{v}\simeq_{(F_{v},\delta)}\psi_{v}$ any $v\in V_{i}$.
Let us recursively modify the unitary elements $\{\omega_{i}\}_{1}^{l}$ in order to construct the global unitary we are looking for. For any $1\leq i\leq l-1$, we fix a subinterval $Y_{i}:=[y_{i,0},y_{i,1}]\subsetneq V_{i}\cap V_{i+1}$ and $\delta>0$ small enough (in particular, we choose $\delta<\epsilon$) in order to be able to invoke  B and obtain a unitary element $\widetilde{w}_{i}$ of $\mathcal{B}$ satisfying the following
$\biggl\{\begin{array}[]{ll}(\widetilde{w}_{i}\phi\widetilde{w}_{i}^{*})_{y}\simeq_{(F_{y},\epsilon)}\psi_{y},\quad\forall y\in Y_{i}\\ \widetilde{w}_{i}(y_{i,0})=w_{i}(y_{i,0})\text{ and }\widetilde{w}_{i}(y_{i,1})=w_{i+1}(y_{i,1})\end{array}$
@@
$\begin{array}[]{ll}w\colon[0,1]\longrightarrow\bigsqcup_{x}B_{x}\\ \hskip 28.45274ptx\longmapsto\biggl\{\begin{array}[]{ll}w_{i}(x)&\text{ if}\quad x\in V_{i}\setminus Y_{i}\\ \widetilde{w}_{i}(x)&\text{ if}\quad x\in Y_{i}\end{array}\end{array}$
from which we conclude that $(w\phi w^{*})_{x}\simeq_{(F_{x},\epsilon)}\psi_{x}$ for any $x\in\bigcup_{1}^{l}V_{i}=[0,1].$
Corollary 3.22.
@@
$w_{n}\phi w_{n}^{*}(a)\underset{n\rightarrow\infty}{\longrightarrow}\psi(a),\quad\forall a\in\mathcal{A}$
@@ C. The classification
Firstly, we say that two continuous field morphisms $\phi,\psi\colon(X,\{A_{x}\}_{x},\mathcal{A})\rightarrow(X,\{B_{x}\}_{x},\mathcal{B})$ are approximately unitarily equivalent, if there exists a sequence $(w_{n})_{n}$ of unitaries of $\mathcal{B}$ such that $w_{n}\phi w_{n}^{*}\underset{n\rightarrow\infty}{\longrightarrow}\psi$ in the point-norm topology. Equivalently, if they are approximately unitarily equivalent as *-homomorphisms.
Secondly, we say that a pair $(X,\{A_{x}\}_{x},\mathcal{A}),(X,\{B_{x}\}_{x},\mathcal{B})]$ of continuous fields satisfies the classification property, termed (CP) for short, if for any scaled $\operatorname{Cu}$-morphism $\alpha\colon\operatorname{Cu}(\mathcal{A})\rightarrow\operatorname{Cu}(\mathcal{B})$ descending to the fibers, there exists a continuous field morphism $\phi\colon\mathcal{A}\rightarrow\mathcal{B}$, unique up to approximate unitary equivalence, such that $\operatorname{Cu}(\phi)=\alpha$. We may write $\mathcal{A},\mathcal{B}]$ has (CP)’ when the context is clear.
@@
Proposition 3.24.
Let $(X,\{(A_{i})_{x}\}_{x},\mathcal{A}_{i})_{i}$ be a family of continuous fields indexed over an upward-directed set. Consider any inductive system $((X,\{(A_{i})_{x}\}_{x},\mathcal{A}_{i}),\phi_{i,j})_{i}$ in $\mathcal{F}_{C^{*}}$. We have the following.
(i) $\mathcal{F}_{C^{*}}\text{-}\lim(\mathcal{A}_{i},\phi_{i,j})\simeq C^{*}\text{-}\lim(\mathcal{A}_{i},\phi_{i,j})$.
(ii) $\mathcal{F}_{C^{*}}\text{-}\lim(\mathcal{A}_{i},\phi_{i,j}),\mathcal{B}]$ has (CP) whenever $\mathcal{A}_{i},\mathcal{B}]$ has (CP) for all $i\in I$.
@@
(i) Consider the inductive system in the category $C^{*}$ and denote by $(\mathcal{A},\phi_{i,\infty})$ its $C^{*}$-limit. For any $x\in X$, consider the inductive system obtained by the projection of the objects and morphisms at the fiber $x$. We obtain another inductive system in $C^{*}$ and we denote its limit by $(A_{x},(\phi_{i,\infty})_{x})$. A standard one-sided intertwining argument shows that there exists a unique *-homomorphism $\pi_{x}\colon\mathcal{A}\rightarrow A_{x}$ compatible with the squares of the intertwining. A fortiori, $(X,\{A_{x}\}_{x},\mathcal{A})$ is a well-defined continuous field over $X$ and $\phi_{i,\infty}$ is a continuous field morphism for any $i\in I$. It is now readily checked that $(\mathcal{A},\phi_{i,\infty})$ satisfies the universal property of inductive limits in the category $\mathcal{F}_{C^{*}}$.
(ii) This is done similarly as in the proof of [13, Proposition 5.2 (iv)]. ∎
In what follows, $\mathcal{C}$ denotes the class of continuous fields that are realized as $\underset{\rightarrow}{\lim}(C[0,1]\otimes A_{i},\phi_{i,i+1})_{i\in\mathbb{N}}$, where $A_{i}$ are unital interval algebras and $\phi_{i,i+1}$ are unital continuous fields morphisms.
@@
Combine  A with  B and the fact that $\mathcal{A},\mathcal{B}]$ has (CP). ∎
Corollary 3.26.
Let $\mathcal{A},\mathcal{B}\in\mathcal{C}$. Then any $\operatorname{Cu}$-isomorphism $\alpha\colon\operatorname{Cu}(A)\simeq\operatorname{Cu}(B)$ descending to the fibers and such that $\alpha([1_{A}])=[1_{B}]$, lifts to a unique (up to approximate unitary equivalence) continuous field isomorphism $\mathcal{A}\simeq\mathcal{B}$.
@@
Remark 3.27 (Computational use of the classification).
The Cuntz semigroup of a $C^{*}$-algebra (and a fortiori of a continuous field) is often very hard to compute. Some results in this direction can be found e.g., in [7, 20, 2]. In our opinion, it seems unlikely that the computation of the Cuntz semigroup of a $\mathcal{C}$-object can be recovered from the one of its fibers, since $\mathrm{K}_{1}$-obstructions will appear in general. (See [2, Theorem 3.4].)
Nevertheless, we believe that the classification theorem obtained could be used to distinguish or relate explicit of continuous fields via approximate intertwining methods similarly as employed in e.g. [9, 10].